OpenStudy (anonymous):

use the fundamental theorem of line integrals to determine if F=(ysinz)i+(xsinz)j+(xycosz)k is conservative

OpenStudy (da_scienceman):

when is a force conservative? is it when the gradient field is zero? can u compute the gradient field? might be m wrong or close to it!

OpenStudy (turingtest):

where are you stuck?

OpenStudy (turingtest):

yes, the fundamental theorem for line integrals states that if the vector field is conservative, the closed line integral is 0

OpenStudy (anonymous):

i found the curl of F to be 0, but i wasnt sure if that theorem was that of the fundamental theorem of line integrals

OpenStudy (anonymous):

If the the curl is 0, it so implies that the vector field is conservative

OpenStudy (turingtest):

right

OpenStudy (turingtest):

no wait not the curl, but the line integral...

OpenStudy (turingtest):

\[\oint\limits_C\nabla f\cdot d\vec r=0\]

OpenStudy (anonymous):

i am confused. does it mean that using the curl to determine if the field is conservative is not consistent with the line integral?

OpenStudy (turingtest):

if the curl is zero that is a conservative vector field, but that doesn't apply the fundamental theorem for line integrals

OpenStudy (turingtest):

so I'm not sure if you solved the problem the way your professor wants you to

OpenStudy (anonymous):

i thought so too. how do i go about it?

OpenStudy (turingtest):

I need to refresh myself on the subject... we need to show that\[\oint\limits_C\vec F\cdot d\vec r=0\]I can think of just picking some closed path and integrating it, but that wouldn't prove the general case holds...

OpenStudy (turingtest):

or at least that\[\int_a^b\vec F\cdot d\vec r\]is path-independent

OpenStudy (turingtest):

the problem seems a bit backwards to me; we already have the vector field, so showing that the curl is zero is the best way to show that it is conservative, but what does that have to do with using the fundamental theorem...?

OpenStudy (anonymous):

beats my mind. i initially thought he wanted us to determine it by showing that a potential exists

OpenStudy (anonymous):

but to show that a potential exists, i first need to find if the field is conservative

OpenStudy (turingtest):

that's not true, we can find the potential function, though it's a pain in the butt again that would not use the fundamental theorem though

OpenStudy (turingtest):

example 3 here shows how to get at the potential function... takes a good minute http://tutorial.math.lamar.edu/Classes/CalcIII/ConservativeVectorField.aspx

OpenStudy (anonymous):

ok. i've been on this for hours. i guess i'll just use curl approach to show that the field is conservative. i cant think of any other approach

OpenStudy (turingtest):

I'll let you know if I get any ideas, good luck!

OpenStudy (anonymous):

thx.

OpenStudy (anonymous):

∮C∇f⋅dr⃗ =0 if the curl is ∇f=0, then their dot product should be equal to a 0. this is how i have decided to go about it

OpenStudy (turingtest):

sounds reasonable to me. The only other way I could think to do it would be to get the potential function and do the line integral along some random path from a to b, and show that it is invariant, but that seems a lot of work

OpenStudy (da_scienceman):

i meant that u might need to compute the first partial derivatives and check if they are equal and conclude if the given function is conservative....what do u think?

OpenStudy (da_scienceman):

theorem: If F = Pi + Qj is continuous in a closed and simply connected region D with first partial derivative satisfying \[\frac{ dP }{ dy}=\frac{ dQ }{ dx }, F = \Pi+Qj\] ...then F is conservative??

OpenStudy (da_scienceman):

ok I dont knw if u can extend this to (i,j,k) as well but for (i,j) m sure it goes that way!

OpenStudy (anonymous):

me too. i actually used a similar system to solve the 2d problems